# Reducing quantum error correction overhead using soft information

### v2

@@ II.3.1 Quantum memory performance
Starting with superconducting qubits, Fig. 2 shows the effect of the measurement classification error probability $p_{\textrm{S}}$ and the two-qubit gate fidelity $\mathcal{F}=1-p$ on the error suppression factor $\Lambda$. As qubit fidelity improves, increasing the code distance has a more significant suppressive effect on the logical error rate, leading to larger $\Lambda$-factors with high $\mathcal{F}$. In the LF-regime, shown in Fig. 2 a), the introduction of soft information to the decoder does not significantly affect $\Lambda$ for any of the physical error rates shown. At $p=0.2\%$ we have a minor $(2.3\pm 0.1)\%$ increase in $\Lambda$ when averaged over the $X$- and $Z$-basis, with the relative gain remaining at $\sim 2\%$ all the way up to $p=0.5\%$. This marginal boost in $\Lambda$ does not significantly improve the physical qubit overhead of an error-corrected device, demonstrating that in a setting where soft measurement errors are sub-dominant ($p_{\textrm{S}}\sim p$), the gain from soft information decoding for superconducting qubits is minimal.
@@
In Fig. 3 we plot $\Lambda$ against the physical qubit fidelity $\mathcal{F}=1-p$ for a surface code memory simulation of neutral atom qubits. For simplicity and ease of comparison with the superconducting qubit simulations, we use square surface codes for these examples despite the high noise bias. As expected, improving qubit fidelities on neutral atoms lead to higher $\Lambda$-values. Due to the strong $Z$-bias and the near-lack of idling errors in the neutral atom noise model, the maximum $\Lambda$ in the LF-regime at $p=0.5\%$ is up to $\Lambda=(7\pm 1)$ for the soft decoder and $\Lambda=(5.8\pm 0.7)$ for the hard decoder. We note that extracting high $\Lambda$-values at low error rates is challenging due to the high number of shots required, leading to large error bars. In the LF-regime, the improvement in $\Lambda$ for soft versus hard decoding ranges from $(16\pm 3)\%$ at $p=0.5\%$ to $(3.0\pm 0.1)\%$ at $p=1\%$, indicating that the biggest gains in error suppression can only be achieved at high qubit fidelity.
Figure 3: LCD on neutral atom qubits: error suppression rate $\Lambda$ versus physical qubit fidelity $\mathcal{F}=1-p$. Data from a simulated quantum memory experiment with $T=10$ rounds of syndrome extraction and $N=10^{6}$ shots per data point. In panel a), we show $\Lambda$ for a noise regime where soft flips are sub-dominant $p_{\textrm{S}}/p=1$, and in panel b) we show $\Lambda$ when soft flips are a significant component of the error model $p_{\textrm{S}}/p=5$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
@@
II.3.2 Measurement time optimisation
@@
For superconducting qubits, we take the SI1000 noise model as our baseline circuit-level noise model and introduce a variation to its error channel probabilities that depends on measurement and gate durations, as described in section IV.1.1. In our parametrisation, the measurement time $\tau_{\textrm{M}}$ alters the probability of depolarising errors while qubits are idling, being measured or reset. The device operation times are set as $\tau_{\textrm{1q}}=20~{}\text{ns}$ and $\tau_{\textrm{2q}}=40~{}\text{ns}$ for one-qubit and two-qubit gates respectively and $\tau_{\textrm{R}}=20~{}\text{ns}$ for resets. The soft measurement response function is likewise dependent on $\tau_{\textrm{M}}$, with a signal-to-noise ratio (SNR) that is proportional to $\tau_{\textrm{M}}$ for measurement times $\tau_{\textrm{M}}\ll\textrm{T}_{1}$. Due to amplitude damping with $\textrm{T}_{1}=100~{}\mu\textrm{s}$, in line with the current state-of-the art devices [1], qubits experience time-dependent decay both while idling and during measurement [37]. For neutral atom qubits, the dominant source of error is the two-qubit error rate—with idle noise being insignificant. We use the same $Z$-biased circuit-level noise model (see section IV.1.2) as done previously, but the measurement response function is varied according to $\tau_{\textrm{M}}$.
[…]
@@
In the neutral atom picture, time-dependent errors during idling are not significant thanks to the long qubit coherence times and fast single-qubit and two-qubit gates [38, 32]. When varying the measurement time, only the classification error probability is affected, leading to both the soft and the hard decoders achieving maximum $\Lambda$ at $\tau_{\textrm{M}}\approx 150~{}\mu\textrm{s}$ as seen in Fig. 4 b). Notably, the improvement in $\Lambda$ for the soft decoder compared to the hard decoder is sustained even for $\tau_{\textrm{M}}>200~{}\mu\textrm{s}$. This is explained by the presence of background noise and decay-related classification errors in the system that persist despite the long measurement time.
II.4 Soft belief propagation decoding
In this section, we investigate whether the advantageous scaling behaviour of soft information decoding seen with LCD can be replicated with higher-accuracy decoders. To study this, we take the belief propagation (BP) pre-decoder and make a set of key modifications that allow it to process soft measurement information, explained in detail in section IV.4. While previous works have tackled the problem of soft information in BP decoders in the limited setting of phenomenological noise [39, 40], our methodology is applicable to general circuit-level noise and does not require additional error nodes. We then test these decoders on surface codes and bivariate bicycle codes under a circuit-level noise simulation, comparing the performance of the soft-information-augmented decoders to their hard-information counterparts.
II.4.1 Soft information in BP-Matching
@@
Our simulations show that the BM variant with soft information added after the hyper-graph generation performs worse than a hard BM decoder. We therefore use the variant with soft information introduced at the belief stage—the soft-belief matching decoder.
@@
Figure 5: BM on superconducting qubits: error suppression rate $\Lambda$ versus physical qubit fidelity $\mathcal{F}=1-p$. Data from a simulated quantum memory experiment with $T=10$ rounds of syndrome extraction and $N=5\times 10^{5}$ shots per data point. In panel a), we show $\Lambda$ for a noise regime where soft flips are sub-dominant $p_{\textrm{S}}/p=1$, and in panel b) we show $\Lambda$ when soft flips are a significant component of the error model $p_{\textrm{S}}/p=5$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
@@
In the LF-regime of Fig. 5 a), we see statistically insignificant $(6\pm 2)\%$- and $(4\pm 0.5)\%$-improvements in $\Lambda$ when $p=0.3\%$ and $p=0.5\%$ respectively. We note that the error bars in this simulation are relatively large due to the large time overhead of the belief propagation algorithm limiting the number of shots taken. In the HF-regime, shown in Fig. 5 b), we see a pattern emerging, with an improvement in $\Lambda$ of to $(16\pm 4)\%$ when $p=0.3\%$ and to $(10\pm 1)\%$ when $p=0.5\%$. Again, we observe that the logical error suppression obtained from soft decoding is more robust to increased classification errors than hard decoding. Compared to the results seen with LCD in section II.3.1, soft BM achieves higher $\Lambda$-values at comparable error rates—reaching $\Lambda=(2.4\pm 0.5)$ at $p=0.3\%$ in the HF-regime versus $\Lambda=(1.78\pm 0.03)$ for soft LCD. Similarly to LCD, soft information does not provide an advantage to BM in the LF-regime where $p_{\textrm{S}}\sim p$, but the performance uplift is significant when soft measurement errors dominate.
II.4.2 Soft information in BP+OSD
@@
The Gross code $[[n,k,d]]\!\!=\!\![[144,12,12]]]$ is among the most promising for near-term demonstrations thanks to its high encoding rate. In this section, we analyse the performance of BB codes under a circuit-level noise model for neutral atom qubits as well as the superconducting qubits, decoding the resulting soft measurements with a BP decoder augmented with soft information. To break the degeneracy problem otherwise inherent in the BP algorithm, we combine soft-BP with Ordered Statistics Decoding (OSD) [43, 44, 41] as a post-processing step. We note that significantly faster decoders have been proposed [45, 46], however these were not available to us when these simulations were conducted.
@@
To evaluate error suppression performance, we take the syndrome extraction circuits in [47] for two instances of BB codes $[[n,k,d]]\!\!=\!\![[72,12,6]]$ and $[[144,12,12]]$, and replace the noise in these circuits with an architecture-specific circuit-level noise model. For the neutral atom simulations, we use a $Z$-biased noise model as described in section IV.1.2, while the superconducting simulations have a SI1000 noise model applied, as detailed in section IV.1.1. As before, we evaluate decoder performance under two noise regimes—LF with $p_{\textrm{S}}=p$ and HF with $p_{\textrm{S}}=5p$, decoding with both a soft-information-aware BP+OSD decoder and a traditional hard BP+OSD decoder. We set up BP with $100$ BP iterations, min-sum as the BP method of choice, combination sweep as the OSD method and an OSD order of 00. The measurement error probabilities obtained from soft information are used to dynamically update the error channel probabilities of the BP+OSD decoder using the update channel probabilities function in [48]. As our figure of merit we use the logical error probability ($P_{\textrm{L}}$) instead of $\Lambda$, as we only used two codes with $d=6$ and $d=12$ for this simulation.
@@
In Fig. 6, we evaluate decoding accuracy for BP+OSD under the LF and HF soft-measurement flip regimes for both superconducting and neutral atom qubit platforms. We find that in the LF-regime, soft information provides more than a $90\%$ reduction in the logical error probability across the physical error rate regime examined compared to hard BP+OSD. In the HF-regime, soft information provides more than a $95\%$ reduction in logical errors compared to hard BP+OSD. Our results clearly show that soft information in BP+OSD significantly enhances the error suppression ability of BB codes, especially at the high physical error rates likely to be seen in near-term devices.
@@
Figure 6: Logical error probability $P_{\textrm{L}}$ as a function of the physical error rate $p$ for bivariate bicycle (BB) codes decoded with BP+OSD. We consider two instances of BB codes: $d\!\!=\!\!6$ for the $[[72,12,6]]$ code and $d\!\!=\!\!12$ for the $[[144,12,12]]$ Gross code. In panels a) and b) we plot the logical error probability $P_{\textrm{L}}$ for simulated neutral atom qubits under two different regimes of soft measurement error probability: $p_{\textrm{S}}/p=1$ for sub-dominant soft measurement errors, and $p_{\textrm{S}}/p=5$ for dominant soft measurement errors. Each data point in panels a) and b) uses $5\times 10^{4}$ shots. In panels c) and d) we plot the logical error probability $P_{\textrm{L}}$ for simulated superconducting qubits using $5\times 10^{5}$ shots per data point, with $p_{\textrm{S}}/p=1$ and $p_{\textrm{S}}/p=5$ in panels c) and d) respectively. Error bars in all the plots correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
III Discussion

### v3

@@ III.1.1 Quantum memory performance
Starting with superconducting qubits, Fig. 2 shows the effect of the measurement classification error probability $p_{\textrm{S}}$ and the two-qubit gate fidelity $\mathcal{F}=1-p$ on the error suppression factor $\Lambda$. For these simulations, we took $10^{6}$ shots per data point. As qubit fidelity improves, increasing the code distance has a more significant suppressive effect on the logical error rate, leading to larger $\Lambda$-factors with high $\mathcal{F}$. In the LF-regime, shown in Fig. 2 a), the introduction of soft information to the decoder does not significantly affect $\Lambda$ for any of the physical error rates shown. At $p=0.2\%$ we have a minor $(2.3\pm 0.1)\%$ increase in $\Lambda$ when averaged over the $X$- and $Z$-basis, with the relative gain remaining at $\sim 2\%$ all the way up to $p=0.5\%$. This marginal boost in $\Lambda$ does not significantly improve the physical qubit overhead of an error-corrected device, demonstrating that in a setting where soft measurement errors are sub-dominant ($p_{\textrm{S}}\sim p$), the gain from soft information decoding for superconducting qubits is minimal.
@@
In Fig. 3 we plot $\Lambda$ against the physical qubit fidelity $\mathcal{F}=1-p$ for a surface code memory simulation of neutral atom qubits, taking $10^{7}$ shots per data point in the LF-regime and $10^{6}$ shots per data point in the HF-regime. For simplicity and ease of comparison with the superconducting qubit simulations, we use square surface codes for these examples despite the high noise bias. As expected, improving qubit fidelities on neutral atoms lead to higher $\Lambda$-values. Due to the strong $Z$-bias and the near-lack of idling errors in the neutral atom noise model, the maximum $\Lambda$ in the LF-regime at $p=0.5\%$ is up to $\Lambda=(7.0\pm 0.4)$ for the soft decoder and $\Lambda=(6.0\pm 0.3)$ for the hard decoder. We note that extracting high $\Lambda$-values at low error rates is challenging due to the high number of shots required, leading to larger error bars. In the LF-regime, the improvement in $\Lambda$ for soft versus hard decoding ranges from $(16\pm 1)\%$ at $p=0.5\%$ to $(4.4\pm 0.1)\%$ at $p=1\%$, indicating that the biggest gains in error suppression can only be achieved at high qubit fidelity.
Figure 3: LCD on neutral atom qubits: error suppression rate $\Lambda$ versus physical qubit fidelity $\mathcal{F}=1-p$. Data from a simulated quantum memory experiment with $T=10$ rounds of syndrome extraction and $N=10^{7}$ shots per data point in panel a) and $N=10^{6}$ in panel b). In panel a), we show $\Lambda$ for a noise regime where soft flips are sub-dominant $p_{\textrm{S}}/p=1$, and in panel b) we show $\Lambda$ when soft flips are a significant component of the error model $p_{\textrm{S}}/p=5$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
@@
Given the differences in $\Lambda$-factors seen for the soft and hard decoders on both platforms, we evaluate the physical qubit footprints that are required to achieve the KiloQuop and MegaQuop regimes using a rotated planar surface code. Shown in table 2, we look at the low-noise regimes of $p=0.2\%$ and $p=0.3\%$ for superconducting qubits and $p=0.5\%$ for neutral atom qubits. We observe that in the LF-regime there are only minor differences in qubit count for soft versus hard decoders. However, in the HF-regime we see a clear advantage for soft decoding in terms of physical qubit counts for both superconducting and neutral atom platforms. In the superconducting case, for $p=0.2\%$ a system decoded using soft LCD can reach the KiloQuop regime with $N=241$ physical qubits, as opposed to $N=337$ qubits that are needed with hard information. In the MegaQuop regime, a soft-decoding system needs $1457$ physical qubits, $13\%$ fewer than the $1681$ needed by a hard decoder. For neutral atoms we find a similarly strong reduction in footprint, particularly in the MegaQuop regime, which takes $N=1457$ qubits for a soft decoder in the HF-regime at $p=0.5\%$ compared to a hard decoder needing $N=2177$ physical qubits, a $33\%$ reduction. These improvements demonstrate that soft information can be a powerful tool for reducing the cost and complexity of future large-scale error-corrected quantum devices.
NA $p$ $0.5\%$ $(6.0\pm 0.3)$ $(7.0\pm 0.4)$ $(16\pm 1)\%$
NA $p$ $1\%$ $(2.89\pm 0.03)$ $(3.01\pm 0.03)$ $(4.4\pm 0.1)\%$
NA $5p$ $0.5\%$ $(2.2\pm 0.1)$ $(2.7\pm 0.2)$ $(20\pm 2)\%$
NA $5p$ $1\%$ $(1.26\pm 0.02)$ $(1.44\pm 0.03)$ $(14.6\pm 0.3)\%$
SC $p$ $0.2\%$ $(2.82\pm 0.09)$ $(2.9\pm 0.1)$ $(2.3\pm 0.1)\%$
SC $p$ $0.3\%$ $(1.86\pm 0.03)$ $(1.90\pm 0.03)$ $(2.3\pm 0.1)\%$
SC $p$ $0.5\%$ $(1.18\pm 0.01)$ $(1.21\pm 0.01)$ $(2.0\pm 0.2)\%$
SC $5p$ $0.2\%$ $(2.52\pm 0.07)$ $(2.78\pm 0.09)$ $(10.6\pm 0.4)\%$
SC $5p$ $0.3\%$ $(1.61\pm 0.02)$ $(1.78\pm 0.02)$ $(10.4\pm 0.2)\%$
SC $5p$ $0.5\%$ $(1.07\pm 0.01)$ $(1.13\pm 0.01)$ $(6.4\pm 0.1)\%$
Table 1: Performance of soft decoders on simulated quantum memory experiments on superconducting (SC) and neutral atom (NA) platforms. For neutral atom platforms, we use a $Z$-biased noise model with bias factor $100$. Both the neutral atom and the superconducting simulations use circuit-level noise with error rate $p$, as described in section VI.1.2 and section VI.1.1 respectively. We vary the amount of classification error $p_{\textrm{S}}$ in the system and show the average error suppression factor $\Lambda$ for soft and hard variants of the local clustering decoder. Each experiment uses $T=10$ syndrome extraction rounds and is repeated for $N=10^{6}$ shots, except the NA simulation at $p_{\textrm{S}}=p$ where $N=10^{7}$.
Platform Decoder $\quad p_{\textrm{S}}\quad$ $\quad p\quad$ $N^{\textrm{hard}}_{\textrm{KiloQuop}}$ $\quad N^{\textrm{soft}}_{\textrm{KiloQuop}}$ $\quad N^{\textrm{hard}}_{\textrm{MegaQuop}}$ $\quad N^{\textrm{soft}}_{\textrm{MegaQuop}}$
[…]
NA LCD $5p$ $0.5\%$ $337~(d_{13})$ $241~(d_{11})$ $2177~(d_{33})$ $1457~(d_{27})$
SC LCD $p$ $0.2\%$ $241~(d_{11})$ $241~(d_{11})$ $1249~(d_{25})$ $1249~(d_{25})$
SC LCD $p$ $0.3\%$ $721~(d_{19})$ $721~(d_{19})$ $4049~(d_{45})$ $3697~(d_{43})$
SC LCD $5p$ $0.2\%$ $337~(d_{13})$ $241~(d_{11})$ $1681~(d_{29})$ $1457~(d_{27})$
SC LCD $5p$ $0.3\%$ $1457~(d_{27})$ $881~(d_{21})$ $6961~(d_{59})$ $4801~(d_{49})$
Table 2: Physical qubit footprint estimates given target logical error rates, based on quantum memory simulations on superconducting (SC) and neutral atom (NA) platforms. The number of physical qubits NN required for a code with distance $d$ (shown in parentheses) to reach a target number of error-free quantum operations (Quops) is estimated based on the rotated planar surface code. For neutral atom platforms we use a $Z$-biased circuit-level noise model with bias factor $100$ detailed in as described in section VI.1.2 and for superconducting qubits we use the circuit-level SI1000 noise model with error rate $p$ as detailed in section VI.1.1. In each simulation, we vary the amount of classification error $p_{\textrm{S}}$ and extract syndromes for $T=10$ rounds, repeating the process for $10^{6}$ shots.
III.1.2 Measurement time optimisation
@@
For superconducting qubits, we take the SI1000 noise model as our baseline circuit-level noise model and introduce a variation to its error channel probabilities that depends on measurement and gate durations, as described in section VI.1.1. In our parametrisation, the measurement time $\tau_{\textrm{M}}$ alters the probability of depolarising errors while qubits are idling, being measured or reset. The device operation times are set as $\tau_{\textrm{1q}}=20~\text{ns}$ and $\tau_{\textrm{2q}}=40~\text{ns}$ for one-qubit and two-qubit gates respectively and $\tau_{\textrm{R}}=20~\text{ns}$ for resets. The soft measurement response function is likewise dependent on $\tau_{\textrm{M}}$, with a signal-to-noise ratio (SNR) that is proportional to $\tau_{\textrm{M}}$ for measurement times $\tau_{\textrm{M}}\ll\textrm{T}_{1}$. Due to amplitude damping with $\textrm{T}_{1}=100~\mu\textrm{s}$, in line with the current state-of-the art devices [1], qubits experience time-dependent decay both while idling and during measurement [46]. For neutral atom qubits, the dominant source of error is the two-qubit error rate—with idle noise being insignificant. We use the same $Z$-biased circuit-level noise model (see section VI.1.2) as done previously, but the measurement response function is varied according to $\tau_{\textrm{M}}$.
In Fig. 4, we vary the code distance $d=\{5,7,9\}$ of rotated planar code quantum memories and plot the error suppression rate $\Lambda$ against the measurement time $\tau_{\textrm{M}}$ for superconducting qubits Fig. 4a) and for neutral atom qubits Fig. 4b). We vary the superconducting qubit measurement time on the interval $\tau_{\textrm{M}}\in[200~\text{ns},$ $1500~\text{ns}]$, and the neutral atom measurement time on the interval $\tau_{\textrm{M}}\in[50~\mu\text{s},$ $300~\mu\text{s}]$. Each data point is calculated from $10^{6}$ shots. On both platforms, we see that for short measurement times the rate of error suppression is decreased. This is explained by the more frequent classification errors that result from a noisier signal. The soft decoders show considerably stronger error suppression than their hard counterparts for small $\tau_{\textrm{M}}$, with average $X$-and $Z$-basis $\Lambda$-factor improved by $(44\pm 1)\%$ for superconducting qubits at $\tau_{\textrm{M}}=300~\text{ns}$ and by $(60\pm 2)\%$ for neutral atom qubits at $\tau_{\textrm{M}}=70~\mu\text{s}$.
@@
In the neutral atom picture, time-dependent errors during idling are insignificant thanks to the long qubit coherence times and fast single-qubit and two-qubit gates [43, 35]. When varying the measurement time, only the classification error probability is affected, leading to both the soft and the hard decoders achieving maximum $\Lambda$ at $\tau_{\textrm{M}}\approx 150~\mu\textrm{s}$ as seen in Fig. 4 b). Notably, the improvement in $\Lambda$ for the soft decoder compared to the hard decoder is sustained even for $\tau_{\textrm{M}}>200~\mu\textrm{s}$. This is due to soft measurement errors that persist due to bright-to-dark transitions, which are not mitigated by a longer measurement time. In contrast, on superconducting qubits the measurement classification fidelity improves with long measurement times as long as $\tau_{\textrm{M}}\ll\textrm{T}_{1}$, leading to convergent $\Lambda$-factors for soft and hard decoding when measurement duration is increased.
III.2 Soft belief propagation decoding
In this section, we investigate whether the advantageous scaling behaviour of soft information decoding seen with LCD can be replicated with higher-accuracy decoders. To study this, we take the belief propagation (BP) pre-decoder and make a set of key modifications that allow it to process soft measurement information, explained in detail in section VI.4. While previous works have tackled the problem of soft information in BP decoders in the limited setting of phenomenological noise [47, 48], our methodology is applicable to general circuit-level noise and does not require additional error nodes. We then test these decoders on surface codes and bivariate bicycle codes under a circuit-level noise simulation, comparing the performance of the soft-information-augmented decoders to their hard-information counterparts.
III.2.1 Soft information in BP-Matching
@@
Our simulations show that the BM variant with soft information added after the hyper-graph generation performs worse than a BM decoder in which soft information is added to at the matching stage. We postulate that if soft information is not included from the beginning, BP may converge to an inaccurate estimate of posterior probabilities, the damage from which cannot be fully recovered by the soft updates to MWPM. We therefore use the variant with soft information introduced at the belief stage—the soft-belief matching decoder.
@@
Figure 5: BM on superconducting qubits: error suppression rate $\Lambda$ versus physical qubit fidelity $\mathcal{F}=1-p$. Data from a simulated quantum memory experiment with $T=10$ rounds of syndrome extraction. Sample size is limited to $N=5\times 10^{5}$ due to slow decoder throughput, causing significant uncertainty in $\Lambda$ at low error rates. In panel a), we show $\Lambda$ for a noise regime where soft flips are sub-dominant $p_{\textrm{S}}/p=1$, and in panel b) we show $\Lambda$ when soft flips are a significant component of the error model $p_{\textrm{S}}/p=5$. Error bars in each plot correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
In the LF-regime of Fig. 5 a), we see statistically insignificant $(6\pm 2)\%$- and $(4\pm 0.5)\%$-improvements in $\Lambda$ when $p=0.3\%$ and $p=0.5\%$ respectively. These results were obtained with $5\times 10^{5}$ shots for each data point, resulting in relatively large error bars. In the HF-regime, shown in Fig. 5 b), we see a pattern emerging, with an improvement in $\Lambda$ of to $(16\pm 4)\%$ when $p=0.3\%$ and to $(10\pm 1)\%$ when $p=0.5\%$. Again, we observe that the logical error suppression obtained from soft decoding is more robust to increased classification errors than hard decoding. Compared to the results seen with LCD in section III.1.1, soft BM achieves higher $\Lambda$-values at comparable error rates—reaching $\Lambda=(2.4\pm 0.5)$ at $p=0.3\%$ in the HF-regime versus $\Lambda=(1.78\pm 0.03)$ for soft LCD. Similarly to LCD, soft information does not provide an advantage to BM in the LF-regime where $p_{\textrm{S}}\sim p$, but the performance uplift is significant when soft measurement errors dominate.
The low numbers of samples in this simulation are due to the large decoding time overheads that limit the number of shots taken. There are two main factors that contribute to this. First, the detector error model needs to be updated for each shot with the corresponding soft information. Using circuit-level noise makes the BP part of the algorithm computationally intensive because unlike the matching graph, the circuit-level Tanner graph has a larger number of edges as each qubit might be connected to multiple syndrome nodes. Second, the Ordered Statistics Decoding (OSD) post-processing step used to decode the BP result is highly time-intensive. Overall, even though soft decoding shows a promising uplift in $\Lambda$-factors when $p_{\textrm{S}}=5p$, the slow decoding speed of BP-Matching makes the technique poorly suited to a high-throughput setting such as a superconducting qubit platform.
@@
III.2.2 Soft information in BP+OSD
@@
The Gross code [[n,k,d]]=[[144,12,12]]][[n,k,d]]\!\!=\!\![[144,12,12]]] is among the most promising for near-term demonstrations thanks to its high encoding rate. In this section, we analyse the performance of BB codes under a circuit-level noise model for neutral atom qubits as well as the superconducting qubits, decoding the resulting soft measurements with a BP decoder augmented with soft information. To break the degeneracy problem otherwise inherent in the BP algorithm, we combine soft-BP with Ordered Statistics Decoding (OSD) [51, 52, 49] as a post-processing step. We note that significantly faster decoders have been proposed [53, 54, 20], however these were not available to us when these simulations were conducted.
@@
To evaluate error suppression performance, we take the syndrome extraction circuits in [55] for two instances of BB codes [[n,k,d]]=[[72,12,6]][[n,k,d]]\!\!=\!\![[72,12,6]] and $[[144,12,12]]$, and replace the noise in these circuits with an architecture-specific circuit-level noise model. For the neutral atom simulations, we use a $Z$-biased noise model as described in section VI.1.2, while the superconducting simulations have a SI1000 noise model applied, as detailed in section VI.1.1. Per standard practice we run syndrome extraction circuits for $N_{\textrm{r}}=d$ rounds, and repeat each simulation for a total of $5\times 10^{5}$ shots per data point. In these simulations, we assume qubit connectivity such that all two-qubit gates specified in the syndrome extraction circuits can be realised natively. As before, we evaluate decoder performance under two noise regimes—LF with $p_{\textrm{S}}=p$ and HF with $p_{\textrm{S}}=5p$, decoding with both a soft-information-aware BP+OSD decoder and a traditional hard BP+OSD decoder. We set up BP with $100$ BP iterations, min-sum as the BP method of choice, combination sweep as the OSD method and an OSD order of 0. In this work we use parallel BP schedule. For more details on the methods used within BP we refer the reader to Ref. [51].
The measurement error probabilities obtained from soft information are used to dynamically update the error channel probabilities of the BP+OSD decoder using the update channel probabilities function in [56]. For our figure of merit, we choose the logical error probability $P_{\textrm{L}}$ instead of $\Lambda$ which was used for our surface code simulations. This choice is due to our simulations only involving two BB codes of the same family with distances $d=6$ and $d=12$, which is insufficient for extracting a $\Lambda$-like error suppression metric.
@@
In Fig. 6, we evaluate decoding accuracy for BP+OSD under the LF and HF soft-measurement flip regimes for both superconducting and neutral atom qubit platforms. We find that in the LF-regime, there is an order-of-magnitude improvement in the logical error rate for superconducting qubits at $p=0.6\%$, and the relative gap between the soft and hard decoders remains consistent all the way down to $p=0.3\%$ for both the distance-$6$ and distance-$12$ codes. In the neutral atom case with heavily $Z$-biased noise, soft decoding decreases the logical error rate by $\sim 12\times$ at $p=3\%$ for both the distance-$6$ and distance-$12$ codes. This improvement widens as error rates decrease, and the gap is also distance-dependent: at $p=1\%$ we see a $30\times$ reduction in $p_{\textrm{L}}$ for $d=6$, and an $80\times$ reduction in $p_{\textrm{L}}$ for $d=12$. Our results clearly show that soft information in BP+OSD significantly enhances the error suppression ability of BB codes, especially at the high physical error rates likely to be seen in near-term devices.
@@
Figure 6: Logical error probability $P_{\textrm{L}}$ as a function of the physical error rate $p$ for bivariate bicycle (BB) codes decoded with BP+OSD. We consider two instances of BB codes: d=6d\!\!=\!\!6 for the $[[72,12,6]]$ code and d=12d\!\!=\!\!12 for the $[[144,12,12]]$ Gross code. In panels a) and b) we plot the logical error probability $P_{\textrm{L}}$ for simulated neutral atom qubits under two different regimes of soft measurement error probability: $p_{\textrm{S}}/p=1$ for sub-dominant soft measurement errors, and $p_{\textrm{S}}/p=5$ for dominant soft measurement errors. Each data point in panels a) and b) uses $5\times 10^{4}$ shots. In panels c) and d) we plot the logical error probability $P_{\textrm{L}}$ for simulated superconducting qubits using $5\times 10^{5}$ shots per data point, with $p_{\textrm{S}}/p=1$ and $p_{\textrm{S}}/p=5$ in panels c) and d) respectively. Error bars in all the plots correspond to hypotheses with a likelihood within a factor of 10 of the maximum likelihood hypothesis, given the sampled data.
III.3 Experimental considerations
III.3.1 Real-time decoding
Our simulations suggest that soft information can reduce the measurement times $\tau_{\textrm{M}}$ needed for syndrome extraction while simultaneously improving the logical fidelity. To realise this benefit in practice, one needs a classical decoder that is able to cope with the increased rate of syndrome information and can perform soft-information-based graph re-weighting in real time. The long measurement times of neutral atom platforms make decoding speed a secondary concern, but in the superconducting qubit setting the very fast QEC cycle times of $\sim 1~\mu\textrm{s}$ dictate very strict timing constraints for the classical decoder. While this challenge may seem daunting, there are several optimisations that can be made to the soft decoding workflow originally proposed in [12] that make it potentially feasible in a real-time application, listed below.
First, the soft measurement probability $P(1\mid\mu)$ can be handled with a fixed precision of $N=8$ bits without compromising logical accuracy, as was found by [16]. This bandwidth can be further reduced with the use of measurement compression schemes, where measurements whose probability substantially deviates from the high-confidence regions of 0 or $1$ are allocated $N$ bits, while measurements with a high-likelihood outcome are transmitted as a single bit. This greatly reduces the bandwidth overhead compared to a naive $N$-bit approach. Second, instead of computing the graph modifications that result from the soft measurement samples in a given shot, one can leverage the finite number of possible soft measurement values to pre-compute an edge weight lookup table (LUT). The LUT consists of a map going from a measurement error probability to the new edge weight that needs to be used for the given measurement error mechanism in the decoding graph. Thanks to finite precision, there are only $2^{N}=256$ possible soft measurement error probabilities that the control system needs to transmit, and as there is a 11-to-11 mapping between a measurement error and a graph refinement event, the entire set of graph modifications only needs to be computed once. Storing the soft-measurement graph refinements in a LUT therefore removes the need for complex arithmetic operations during runtime and paves the way for low-latency soft decoding implementations in dedicated hardware.
In the case of decoding bivariate bicycle codes, recent developments in fast BP-based decoders such as Ref. [20] offer hope that real-time soft information decoding of high-encoding rate qLDPC codes may become possible in the near future. Here too, tracking the soft measurements with finite precision and pre-computing all the necessary hyper-edge modifications on the Tanner graph subject to soft measurement errors may prove essential for a high-throughput implementation.
III.3.2 Measurement model fine-tuning
[…]
On neutral atom platforms, the properties of the imaging pipeline may cause deviations between our model (cf. [Section VI.A.2]) and a real-world measurement response. Optical effects such as vignetting of the image near the edges of the array, brightness crosstalk from neighbouring qubits, or other lens aberrations can decrease the signal-to-noise ratio. Additionally, sensor-based noise sources such as dark current non-uniformity and thermal fluctuations can lead to increased background noise [Radnaev et al. (2024)]. These effects can be mitigated by fitting unique per-qubit probability density functions based on a reference set of measurement values, with each PDFs incorporating optical and sensor-based noise components.
IV Discussion